OpenStudy (anonymous):

The total revenue R earned (in dollars) from producing a gift box of candles is given by R(p)=-10p^2+800p, where p is the price per unit. a. Find the revenues when the prices per box are $20, $25, and $30 b. Find the unit price that will yield a maximum revenue. What is the maximum revenue? Explain your results.

OpenStudy (anonymous):

Part a is simply plug and chug stuff that you should use a calculator to solve. Part b, however, requires either a graph of the equation or its derivative. If you're looking to use the latter, set it (-20p+800) equal to zero to solve for the price that maximizes revenue.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I got part a, but when i graphed part b I got a straight line and not a parabola. do you have any idea why I'm getting that?

OpenStudy (anonymous):

Did you graph the equation I gave you or the one you had to start?

OpenStudy (anonymous):

The derivative should work for b -20p+800=0 -20p=-800 p=40

OpenStudy (anonymous):

I don't mean to overwhelm you, but, since this problem is a parabola facing downwards, the max is the vertex; so you could find your max by completing the square as well. Would that be easier?

OpenStudy (anonymous):

ok that would be easier thank you I got it now!

OpenStudy (anonymous):

yw ;)